OpenStudy (anonymous):

Integrate the following by expanding the partial fractions: ∫[cos(x)]/[sin^2(x)-sin(x)-2] dx

OpenStudy (anonymous):

\[\frac{\cos (x)}{(\sin (x)-2) (\sin (x)+1)}\]

OpenStudy (anonymous):

cos(x)/[(sin(x)-2)(sin(x)+1)] A B ---- + ---- sin(x)-2 sin(x)+1 A=(1/3)cos(x) B=-(1/3)cos(x)

OpenStudy (anonymous):

Let u = sinx therefore du = cosx dx Substituting these two into your integral makes it easier to solve so now you have: \[\int\limits_{}^{} du / (u + 1)(u - 2)du\]

OpenStudy (anonymous):

so u need to integrate: \[(1/3)\int\limits_{}^{}\cos(x)/(\sin(x)-2)-\cos(x)/(\sin(x)+1)\]

OpenStudy (anonymous):

1 / (u +1)(u - 2) = A / (u + 1) + B / (u - 2)

OpenStudy (anonymous):

Multiply both sides by the lowest common denominator which is (u+ 1)(u - 2)

OpenStudy (anonymous):

Therefore: 1 = A(u - 2) + B(u + 1) If u = 2 ; 1 = 3B, B = 1/3 if u = -1; 1 = -3A, A = -1/3

OpenStudy (anonymous):

So now you have\[1/3\int\limits_{}^{} 1 / [ (u - 2) - 1 / (u + 1) ] du\]

OpenStudy (anonymous):

1/3 [ ln|u - 2| - ln | u + 2| ] + C

OpenStudy (anonymous):

Sorry previous post should be u + 1 1/3 ln|(u - 2)/(u + 1)|

OpenStudy (anonymous):

NICE same as mine tutorial answer thanks =)

OpenStudy (anonymous):

you're welcome! So you get it?

OpenStudy (anonymous):

Then just back substitute for u = sinx

OpenStudy (anonymous):

yea got the idea on it now working it out on my paper =)

OpenStudy (anonymous):

cool!